# Parrondo’s paradox in susceptible-infectious-susceptible dynamics over periodic temporal networks

### v1

@@ 3.5 Perturbation theory for the largest Floquet exponent
Regardless of the presence or absence of the Parrondo’s paradox as one varies $\beta$, Fig. 2 suggests that the relationship between the largest Floquet exponent (i.e., $\lambda_{\text{F}}$) and $\beta$ is apparently close to linear for an extended range of $\beta$. Therefore, we derive the first-order approximation of $\lambda_{\text{F}}$ in terms of $\beta$ and assess its accuracy in this section.
@@
Now we validate our first-order approximation for $2\times 2$ periodic switching networks with $\ell=2$. We compare $\lambda_{\text{F}}$ between the exact value and linear approximation in Fig. 6 for eight arbitrarily selected two-node periodic switching networks. Figures 6(a)–(d) are for four networks showing the Parrondo’s paradox. We find that the first-order approximation is sufficiently accurate for a range of $\beta$ including the range in which the paradox occurs. The first-order approximation is also accurate for four periodic switching networks in which the paradox is absent in the entire range of $\beta$ (see Fig. 6(e)–(h)). Figure 6 suggests that the first-order approximation tends to be more accurate when the paradox is absent (see Fig. 6(e)–(h)) than present (see Fig. 6(a)–(d)).
Figure 6: First-order approximation to the largest Floquet exponent for eight two-node periodic switching networks with $k=2$, $T=7$, and $\tau_{1}=5$. We set $\mu=0.5$. (a) $\textbf{A}^{(1)}=\begin{pmatrix}17&5\\ 5&0\end{pmatrix}$ and $\textbf{A}^{(2)}=\begin{pmatrix}0&5\\ 5&15\end{pmatrix}$. (b) $\textbf{A}^{(1)}=\begin{pmatrix}12&5\\ 5&1\end{pmatrix}$ and $\textbf{A}^{(2)}=\begin{pmatrix}1&5\\ 5&15\end{pmatrix}$. (c) $\textbf{A}^{(1)}=\begin{pmatrix}17&0\\ 0&0\end{pmatrix}$ and $\textbf{A}^{(2)}=\begin{pmatrix}0&0\\ 0&15\end{pmatrix}$. (d) $\textbf{A}^{(1)}=\begin{pmatrix}25&2\\ 3&4\end{pmatrix}$ and $\textbf{A}^{(2)}=\begin{pmatrix}6&1\\ 7&30\end{pmatrix}$. (e) $\textbf{A}^{(1)}=\begin{pmatrix}3&2\\ 2&3\end{pmatrix}$ and $\textbf{A}^{(2)}=\begin{pmatrix}4&8\\ 6&26\end{pmatrix}$. (f) $\textbf{A}^{(1)}=\begin{pmatrix}10&5\\ 2&15\end{pmatrix}$ and $\textbf{A}^{(2)}=\begin{pmatrix}4&3\\ 6&8\end{pmatrix}$. (g) $\textbf{A}^{(1)}=\begin{pmatrix}2&4\\ 4&2\end{pmatrix}$ and $\textbf{A}^{(2)}=\begin{pmatrix}3&8\\ 8&3\end{pmatrix}$. (h) $\textbf{A}^{(1)}=\begin{pmatrix}30&5\\ 10&7\end{pmatrix}$ and $\textbf{A}^{(2)}=\begin{pmatrix}3&6\\ 6&3\end{pmatrix}$.
As an application of the first-order approximation of $\lambda_{\text{F}}$, we examine the fraction of randomly generated periodic switching networks for which the Parrondo’s paradox occurs, as we did in section 3.3. Note that the epidemic threshold is given in the form of $\mu/\lambda_{\max}$ for the periodic switching network under the first-order approximation as well as for the static networks. Therefore, if the first-order approximation is accurate, the Parrondo’s paradox happens if $\lambda_{\max}(\overline{\textbf{A}})$ is larger than both $\lambda_{\max}(\textbf{A}^{(1)})$ and $\lambda_{\max}(\textbf{A}^{(2)})$ or smaller than both of them. In the same manner as that in section 3.3, we generated $10^{4}$ pairs of $\textbf{A}^{(1)}$ and $\textbf{A}^{(2)}$ and computed $\lambda_{\max}(\overline{\textbf{A}})$, $\lambda_{\max}(\textbf{A}^{(1)})$, and $\lambda_{\max}(\textbf{A}^{(2)})$. We found that $\lambda_{\max}(\overline{\textbf{A}})$ is smaller than both $\lambda_{\max}(\textbf{A}^{(1)})$ and $\lambda_{\max}(\textbf{A}^{(2)})$ for $31.7\%$ pairs of $\textbf{A}^{(1)}$ and $\textbf{A}^{(2)}$. There was no case in which $\lambda_{\max}(\overline{\textbf{A}})$ is larger than $\lambda_{\max}(\textbf{A}^{(1)})$ and $\lambda_{\max}(\textbf{A}^{(2)})$. These results are similar to those obtained for $\lambda_{\text{F}}$ (see section 3.3).
@@ 4 Discussion
We showed a Parrondo’s paradox in the SIS dynamics on periodically switching temporal networks in which an epidemic decays over a long time while the dynamics is super-critical (i.e., above the epidemic threshold) for the momentary static network at any point of time. We primarily showed the paradox in simple cases involving a periodic alternation of two $2\times 2$ adjacency matrices. Larger networks (with $N=2000$ nodes) with two subpopulations generated by the SBM also showed the paradox, even for a single run of stochastic SIS dynamics. We further verified that the paradox occurred with adjacency matrices of larger sizes, up to five nodes. However, we found that the fraction of random networks that show the Parrondo’s paradox decreased as the number of nodes increases. This is an important limitation of the present study. Further investigating the generality of the Parrondo’s paradox in terms of the number of nodes or subpopulations, the number of static networks constituting a periodic switching network, different epidemic process models, and different network models such as the metapopulation model and hypergraphs, warrants future work.
@@
We pointed out the relationship between the growth rate of the epidemics (including the epidemic threshold, at which the growth rate is equal to 00) and the Floquet exponent. Leveraging Floquet theory to characterize and intervene into epidemic processes on periodic temporal networks may be fruitful. For example, the periodic matrix obtained by the Floquet decomposition (i.e., matrix $\textbf{P}(t)$ in Appendix A) may help us to quantify the amplitude of antiphase oscillation between $x_{i}(t)$ and $x_{j}(t)$ (with $j\neq i$) or change it through manipulation of $\textbf{P}(t)$.
@@
In the context of COVID-19, effects on mitigating spread of infection were numerically simulated on physical proximity and face-to-face contact data and compared between “rotating” and “on-off” strategies with two-day or biweekly periodic schedules [86]. The rotating strategy by definition imposes that half the population is allowed to go out in odd days or odd weeks, and the other half is allowed to go out in even days or even weeks. In contrast, the on-off strategy allows the entire population to go for work in one day (or one week) and not on the next day (or week), and such a two-day or biweekly pattern is repeated. As expected, any of these four strategies helped towards decreasing the basic reproduction number below $1$. Other studies also investigated effects of different on-off strategies on suppressing epidemics [87, 88]. Additionally, it was shown in Ref. [86] that among the four strategies, the best and worst strategies were the rotating strategy with the biweekly cycle and the on-off strategy with two-day cycle, respectively. Superiority of a rotating strategy to an on-off strategy was also shown in a different study [89]. These comparative results suggest that different quarantine or curfew strategies based on periodic switching networks are feasible and that their efficiency depends on the periodic switching network. Investigating the possibility of the Parrondo’s paradox and its possible roles in suppressing population-wide infection in these and other practical disease spreading models warrants future study. In particular, both these previous studies and the present work suggest that eliciting anti-phase oscillations between different communities might be a key to suppressing infection while keeping the overall activity level of individuals relatively high.
Appendix Appendix A Floquet theory
@@
Definition A.1 (Periodic system).
@@
Theorem A.2 implies that the solution of (32) is represented as
@@
Appendix Appendix B Simulations of the stochastic SIS model and nonlinear IBA
@@
We show in Fig. 7(a)–(d) the time course of $x_{1}(t)$ and $x_{2}(t)$ in a single run of stochastic SIS dynamics simulated using the Gillespie algorithm. We employed the four values of $\beta$ used in Fig. 2(b)–(e). The fluctuations in $x_{1}(t)$ and $x_{2}(t)$ are large because the results shown are from a single run of stochastic SIS dynamics. However, there is some indication of the Parrondo’s paradox and anti-phase oscillation in Fig. 7(b) and (c). For a value of $\beta$ at which the paradox occurs according to our theoretical analysis (i.e., $\beta=0.032$; shown in Fig. 7(c)), we further carried out a single run of stochastic simulation on each of the two static networks with $N=2000$ nodes constituting the periodic switching network used in Fig. 7. Figure 8(a) and (b) represents an example time course of $x_{1}(t)$ and $x_{2}(t)$ on the first and second static network, respectively. We find that $x_{1}(t)$ and $x_{2}(t)$ do not decay exponentially over time in each static network in this run. Therefore, although not all the runs show this behavior, the Parrondo’s paradox can occur in a single run of stochastic simulation.
@@
We then simulate the same stochastic SIS dynamics 100 times for each value of $\beta$ and average $x_{1}(t)$ and $x_{2}(t)$ over the 100 runs. We show the results, together with the standard deviation of $x_{1}(t)$ and $x_{2}(t)$, in Fig. 7(e)–(h). We find a consistent tendency of the Parrondo’s paradox and anti-phase oscillation in the averaged time course of $x_{1}(t)$ and $x_{2}(t)$ (see Fig. 7(f) and (g)).
@@
For comparison, we also show numerically simulated solutions of the nonlinear IBA, Eq. (2), and its linearized variant, Eq. (3), for the $2\times 2$ periodic switching network in the same figure. The time courses of $x_{1}(t)$ and $x_{2}(t)$ for the linearized IBA shown in Fig. 7(e)–(h) are the same as those shown in Fig. 2(b)–(e). We observe that the nonlinear IBA shows less pronounced Parrondo’s paradox and that the average time courses of the stochastic simulations when the paradox occurs are close to those for the nonlinear IBA (see Fig. 7(f) and (g)). We conclude that stochastic agent-based simulations also show the Parrondo’s paradox, albeit to a lesser extent than the linearized IBA with a small number of communities shown in the main text.
Appendix Appendix C Anti-phase oscillatory behavior with examples
@@
We measure $q$, i.e., the extent of the anti-phase oscillation defined in section 3.4, for the eight two-node periodic switching networks that we used in Fig. 6. The upper subpanel in each of Fig. 9(a)–(h) shows $\lambda_{\text{F}}$, $\lambda_{\max}(\textbf{M}^{(1)})$, and $\lambda_{\max}(\textbf{M}^{(2)})$, and the range of $\beta$ in which the paradox occurs, similar to Fig. 2. We find that the first four networks showing the Parrondo’s paradox (i.e., Fig. 9(a)–(d)) produce large $q$ values (i.e., above approximately $0.8$) when $\beta$ is close to the epidemic threshold. The largest $q$ value attained in these four networks is $1$, indicating perfect anti-phase dynamics between $x_{1}(t)$ and $x_{2}(t)$. The range of $\beta$ in which $q=1$ roughly coincides with where the Parrondo’s paradox occurs. However, in one of the four periodic switching networks, shown in Fig. 9(c), the range of $\beta$ in which $q=1$ extends to larger values of $\beta$ far beyond where the paradox occurs. In contrast, for the other four networks that do not show the paradox (Fig. 9(e)–(h)), the largest $q$ value is substantially smaller than for the first four networks, while $q$ values are peaked near the epidemic thresholds for each momentarily static network, except in Fig. 9(g). Note that $q=0.5$ implies that the anti-phase behavior exists only half of the time so one cannot really regard this as anti-phase behavior. In Fig. 9(g), we find $q=0$, which is the complete absence of anti-phase dynamics, for all values of $\beta$.
@@
Figure 9: Fraction of the time showing anti-phase behavior, $q$, in eight two-node periodic switching networks with $\ell=2$, $T=7$, and $\tau_{1}=5$. The eight networks used are the same as those used in Fig. 6. We recall that the Parrondo’s paradox behavior occurs for the first four networks in a range of $\beta$ values, i.e., shaded regions, and it does not occur for the last four networks. We compare $q$ with $\lambda_{\text{F}}$, $\lambda_{\max}(\textbf{M}^{(1)})$, and $\lambda_{\max}(\textbf{M}^{(2)})$ as a function of the infection rate, $\beta$, for each periodic switching network. We set $\mu=0.5$. The lines for $\lambda_{\text{F}}$, $\lambda_{\max}(\textbf{M}^{(1)})$, and $\lambda_{\max}(\textbf{M}^{(2)})$ are the same as those shown in Fig. 6. Observe that there is a strong association between the Parrondo’s paradox behavior and the appearance of large $q$ values.

### v2

@@ 3.5 Perturbation theory for the largest Floquet exponent
Regardless of the presence or absence of the Parrondo paradox as one varies $\beta$, Fig. 3 suggests that the relationship between the largest Floquet exponent (i.e., $\lambda_{\text{F}}$) and $\beta$ is apparently close to linear for an extended range of $\beta$. Therefore, we derive the first-order approximation of $\lambda_{\text{F}}$ in terms of $\beta$ and assess its accuracy in this section.
@@
Now we validate our first-order approximation for $2\times 2$ periodic switching networks with $\ell=2$. We compare $\lambda_{\text{F}}$ between the exact value and linear approximation in Fig. 7 for eight arbitrarily selected two-node periodic switching networks. Figures 7(a)–(d) are for four networks showing the Parrondo paradox. We find that the first-order approximation is sufficiently accurate for a range of $\beta$ including the range in which the paradox occurs. The first-order approximation is also accurate for four periodic switching networks in which the paradox is absent in the entire range of $\beta$ (see Fig. 7(e)–(h)). Figure 7 suggests that the first-order approximation tends to be more accurate when the paradox is absent (see Fig. 7(e)–(h)) than present (see Fig. 7(a)–(d)).
Figure 7: First-order approximation to the largest Floquet exponent for eight two-node periodic switching networks with $k=2$, $T=7$, and $\tau_{1}=5$. We set $\mu=0.5$. We show the largest Floquet exponent (i.e., $\lambda_{F}$) in black, largest eigenvalue of $\textbf{M}^{(1)}$ in blue, that of $\textbf{M}^{(2)}$ in magenta, and the first-order approximation of $\lambda_{F}$ in red.
As an application of the first-order approximation of $\lambda_{\text{F}}$, we examine the fraction of randomly generated periodic switching networks for which the Parrondo paradox occurs, as we did in section 3.3. Note that the epidemic threshold is given in the form of $\mu/\lambda_{\max}$ for the periodic switching network under the first-order approximation as well as for the static networks. Therefore, if the first-order approximation is accurate, the Parrondo paradox happens if $\lambda_{\max}(\overline{\textbf{A}})$ is larger than both $\lambda_{\max}(\textbf{A}^{(1)})$ and $\lambda_{\max}(\textbf{A}^{(2)})$ or smaller than both of them. In the same manner as that in section 3.3, we generated $10^{4}$ pairs of $\textbf{A}^{(1)}$ and $\textbf{A}^{(2)}$ and computed $\lambda_{\max}(\overline{\textbf{A}})$, $\lambda_{\max}(\textbf{A}^{(1)})$, and $\lambda_{\max}(\textbf{A}^{(2)})$. We found that $\lambda_{\max}(\overline{\textbf{A}})$ is smaller than both $\lambda_{\max}(\textbf{A}^{(1)})$ and $\lambda_{\max}(\textbf{A}^{(2)})$ for $31.7\%$ pairs of $\textbf{A}^{(1)}$ and $\textbf{A}^{(2)}$. There was no case in which $\lambda_{\max}(\overline{\textbf{A}})$ is larger than $\lambda_{\max}(\textbf{A}^{(1)})$ and $\lambda_{\max}(\textbf{A}^{(2)})$. These results are similar to those obtained for $\lambda_{\text{F}}$ (see section 3.3).
@@ 4 Discussion
We showed a Parrondo’s paradox in the SIS dynamics on periodically switching temporal networks in which an epidemic decays over a long time while the dynamics is super-critical (i.e., above the epidemic threshold) for the momentary static network at any point of time. We primarily showed the paradox in simple cases involving a periodic alternation of two $2\times 2$ adjacency matrices. Larger networks (with $N=2000$ nodes) with two subpopulations generated by the SBM also showed the paradox, even for a single run of stochastic SIS dynamics. We further verified that the paradox occurred with adjacency matrices of larger sizes, up to five nodes. However, we found that the fraction of random networks that show the Parrondo paradox decreased as the number of nodes increases. This is an important limitation of the present study. Further investigating the generality of the Parrondo paradox in terms of the number of nodes or subpopulations, the number of static networks constituting a periodic switching network, different epidemic process models, and different network models such as the metapopulation model and hypergraphs, warrants future work.
@@
We pointed out the relationship between the growth rate of the epidemics (including the epidemic threshold, at which the growth rate is equal to 00) and the Floquet exponent. Leveraging Floquet theory to characterize and intervene into epidemic processes on periodic temporal networks may be fruitful. For example, the periodic matrix obtained by the Floquet decomposition (i.e., matrix $\textbf{P}(t)$ in Appendix B) may help us to quantify the amplitude of antiphase oscillation between $x_{i}(t)$ and $x_{j}(t)$ (with $j\neq i$) or change it through manipulation of $\textbf{P}(t)$.
@@
In the context of COVID-19, effects on mitigating spread of infection were numerically simulated on physical proximity and face-to-face contact data and compared between “rotating” and “on-off” strategies with two-day or biweekly periodic schedules [97]. The rotating strategy by definition imposes that half the population is allowed to go out in odd days or odd weeks, and the other half is allowed to go out in even days or even weeks. In contrast, the on-off strategy allows the entire population to go for work in one day (or one week) and not on the next day (or week), and such a two-day or biweekly pattern is repeated. As expected, any of these four strategies helped towards decreasing the basic reproduction number below $1$. Other studies also investigated effects of different on-off strategies on suppressing epidemics [98, 99]. Additionally, it was shown in Ref. [97] that among the four strategies, the best and worst strategies were the rotating strategy with the biweekly cycle and the on-off strategy with two-day cycle, respectively. Superiority of a rotating strategy to an on-off strategy was also shown in a different study [100]. These comparative results suggest that different quarantine or curfew strategies based on periodic switching networks are feasible and that their efficiency depends on the periodic switching network. Investigating the possibility of the Parrondo paradox and its possible roles in suppressing population-wide infection in these and other practical disease spreading models warrants future study. In particular, both these previous studies and the present work suggest that eliciting anti-phase oscillations between different communities might be a key to suppressing infection while keeping the overall activity level of individuals relatively high.
Appendix Appendix A Proof that all the entries of $e^{\beta\textbf{E}(T)}$ are positive
Here, we show that each entry of the matrix exponential of the time-aggregated adjacency matrix $\textbf{E}(T)=\int_{0}^{T}\textbf{A}(t)\text{d}t$ is positive under the assumption that $\textbf{E}(T)$ is nonnegative and irreducible (i.e., which corresponds to when it is associated with a strongly connected graph with positive edge weights). Recall that the matrix exponential is defined by $e^{\beta\textbf{E}(T)}=\sum_{k=0}^{\infty}\frac{1}{k!}[\beta\textbf{E}(T)]^{k}$, which has entries
$[e^{\beta\textbf{E}(T)}]_{ij}=\sum_{k=0}^{\infty}\frac{1}{k!}[\beta\textbf{E}(T)]_{ij}^{k}=\sum_{k=0}^{\infty}\frac{1}{k!}\beta^{k}[\textbf{E}(T)^{k}]_{ij}.$ (31)
Because $\textbf{E}(T)$ is nonnegative and irreducible, for any $(i,j)\in\{1,\ldots,N\}\times\{1,\ldots,N\}$, there exists a $k$ such that $\frac{1}{k!}\beta^{k}[{\bf E}(T)^{k}]_{ij}>0$. Specifically, $\frac{1}{k!}\beta^{k}[{\bf E}(T)^{k}]_{ij}$ is strictly positive if there is at least one path from $i$ to $j$ having length $k$. Notably, $[{\bf E}(T)^{k}]_{ij}$ equals the product of edge weights along such a path, which is summed across all possible paths. Let $\overline{k}$ denote such a $k$ value. Then, we note that $\frac{1}{k!}\beta^{k}[{\bf E}(T)^{k}]_{ij}$ must be nonnegative for any $k\neq\overline{k}$, since ${\bf E}(T)$ has nonnegative entries and the product and summation of nonnegative numbers remain nonnegative. Thus, for any $i,j$, the right-hand side of Eq. (31) is a positive number (i.e., the term with $k=\overline{k}$) plus the sum of nonnegative numbers (i.e., the terms with $k\neq\overline{k}$), which is a positive number.
Appendix Appendix B Floquet theory
@@
Definition B.1 (Periodic system).
@@
Theorem B.2 implies that the solution of (33) is represented as
@@
Appendix Appendix C Simulations of the stochastic SIS model and nonlinear IBA
@@
We show in Fig. 8(a)–(d) the time course of $x_{1}(t)$ and $x_{2}(t)$ in a single run of stochastic SIS dynamics simulated using the Gillespie algorithm. We employed the four values of $\beta$ used in Fig. 3(b)–(e). The fluctuations in $x_{1}(t)$ and $x_{2}(t)$ are large because the results shown are from a single run of stochastic SIS dynamics. However, there is some indication of the Parrondo paradox and anti-phase oscillation in Fig. 8(b) and (c). For a value of $\beta$ at which the paradox occurs according to our theoretical analysis (i.e., $\beta=0.032$; shown in Fig. 8(c)), we further carried out a single run of stochastic simulation on each of the two static networks with $N=2000$ nodes constituting the periodic switching network used in Fig. 8. Figure 9(a) and (b) represents an example time course of $x_{1}(t)$ and $x_{2}(t)$ on the first and second static network, respectively. We find that $x_{1}(t)$ and $x_{2}(t)$ do not decay exponentially over time in each static network in this run. Therefore, although not all the runs show this behavior, the Parrondo paradox can occur in a single run of stochastic simulation.
@@
We then simulate the same stochastic SIS dynamics 100 times for each value of $\beta$ and average $x_{1}(t)$ and $x_{2}(t)$ over the 100 runs. We show the results, together with the standard deviation of $x_{1}(t)$ and $x_{2}(t)$, in Fig. 8(e)–(h). We find a consistent tendency of the Parrondo paradox and anti-phase oscillation in the averaged time course of $x_{1}(t)$ and $x_{2}(t)$ (see Fig. 8(f) and (g)).
@@
For comparison, we also show numerically simulated solutions of the nonlinear IBA, Eq. (2), and its linearized variant, Eq. (3), for the $2\times 2$ periodic switching network in the same figure. The time courses of $x_{1}(t)$ and $x_{2}(t)$ for the linearized IBA shown in Fig. 8(e)–(h) are the same as those shown in Fig. 3(b)–(e). We observe that the nonlinear IBA shows less pronounced Parrondo’s paradox and that the average time courses of the stochastic simulations when the paradox occurs are close to those for the nonlinear IBA (see Fig. 8(f) and (g)). We conclude that stochastic agent-based simulations also show the Parrondo paradox, albeit to a lesser extent than the linearized IBA with a small number of communities shown in the main text.
Appendix Appendix D Anti-phase oscillatory behavior with examples
@@
We measure $q$, i.e., the extent of the anti-phase oscillation defined in section 3.4, for the eight two-node periodic switching networks that we used in Fig. 7. The upper subpanel in each of Fig. 10(a)–(h) shows $\lambda_{\text{F}}$, $\lambda_{\max}(\textbf{M}^{(1)})$, and $\lambda_{\max}(\textbf{M}^{(2)})$, and the range of $\beta$ in which the paradox occurs, similar to Fig. 3. We find that the first four networks showing the Parrondo paradox (i.e., Fig. 10(a)–(d)) produce large $q$ values (i.e., above approximately $0.8$) when $\beta$ is close to the epidemic threshold. The largest $q$ value attained in these four networks is $1$, indicating perfect anti-phase dynamics between $x_{1}(t)$ and $x_{2}(t)$. The range of $\beta$ in which $q=1$ roughly coincides with where the Parrondo paradox occurs. However, in one of the four periodic switching networks, shown in Fig. 10(c), the range of $\beta$ in which $q=1$ extends to larger values of $\beta$ far beyond where the paradox occurs. In contrast, for the other four networks that do not show the paradox (Fig. 10(e)–(h)), the largest $q$ value is substantially smaller than for the first four networks, while $q$ values are peaked near the epidemic thresholds for each momentarily static network, except in Fig. 10(g). Note that $q=0.5$ implies that the anti-phase behavior exists only half of the time so one cannot really regard this as anti-phase behavior. In Fig. 10(g), we find $q=0$, which is the complete absence of anti-phase dynamics, for all values of $\beta$.
@@
Figure 10: Fraction of the time showing anti-phase behavior, $q$, in eight two-node periodic switching networks with $\ell=2$, $T=7$, and $\tau_{1}=5$. The eight networks used are the same as those used in Fig. 7. We recall that the Parrondo paradox behavior occurs for the first four networks in a range of $\beta$ values, i.e., shaded regions, and it does not occur for the last four networks. We compare $q$ with $\lambda_{\text{F}}$, $\lambda_{\max}(\textbf{M}^{(1)})$, and $\lambda_{\max}(\textbf{M}^{(2)})$ as a function of the infection rate, $\beta$, for each periodic switching network. We set $\mu=0.5$. The lines for $\lambda_{\text{F}}$, $\lambda_{\max}(\textbf{M}^{(1)})$, and $\lambda_{\max}(\textbf{M}^{(2)})$ are the same as those shown in Fig. 7. Observe that there is a strong association between the Parrondo paradox behavior and the appearance of large $q$ values.